OpenStudy (anonymous):

What is the value of 7P3? 21 150 210 375 What is the value of 4C2? 6 8 16 80 There are 20 books on the summer reading list. In how many different ways can you choose one per week for six weeks? 120 21,400 10,350,200 27,907,200 The probability that a spinner will land on a blue section is 1over4. What is the probability that the spinner will not land on a blue section? 1over4 1over2 3over4 1

OpenStudy (anonymous):

1. \[7 \times 6 \times 5 = 210\] 2. \[\frac{4 \times 3}{2}=6\]

OpenStudy (anonymous):

1. \[7 \times 6 \times 5 = 210\] 2. \[\frac{4 \times 3}{2}=6\] 3. 20P6\[20 \times 19 \times 18 \times 17 \times 16 \times 15\]

OpenStudy (anonymous):

Did you understand?